# DenseDPO: Fine-Grained Temporal Preference Optimization for Video Diffusion Models

### v1

@@ Abstract
Direct Preference Optimization (DPO) has recently been applied as a post-training technique for text-to-video diffusion models. To obtain training data, annotators are asked to provide preferences between two videos generated from independent noise. However, this approach prohibits fine-grained comparisons, and we point out that it biases the annotators towards low-motion clips as they often contain fewer visual artifacts. In this work, we introduce DenseDPO, a method that addresses these shortcomings by making three contributions. First, we create each video pair for DPO by denoising corrupted copies of a ground truth video. This results in aligned pairs with similar motion structures while differing in local details, effectively neutralizing the motion bias. Second, we leverage the resulting temporal alignment to label preferences on short segments rather than entire clips, yielding a denser and more precise learning signal. With only one-third of the labeled data, DenseDPO greatly improves motion generation over vanilla DPO, while matching it in text alignment, visual quality, and temporal consistency. Finally, we show that DenseDPO unlocks automatic preference annotation using off-the-shelf Vision Language Models (VLMs): GPT accurately predicts segment-level preferences similar to task-specifically fine-tuned video reward models, and DenseDPO trained on these labels achieves performance close to using human labels. Additional results are available at https://snap-research.github.io/DenseDPO/.
@@ 2 Related Work
Preference learning for image diffusion. Inspired by the success of human feedback learning in language modeling [50, 3], similar approaches have been adapted to image generation. One major line of work focuses on training reward models from human preference labels [77, 74, 73, 89, 33], these models can be used as loss functions to optimize generators by direct gradient backpropagation [13, 75, 9, 31, 57, 77] or policy gradients [5, 18]. Another line of work uses predicted rewards on training data to re-weight the diffusion loss [38], or only trains on high-scoring samples [16, 14, 39]. However, all these methods require an explicitly trained reward model, and may suffer from the reward hacking issue [77, 13]. In contrast, Diffusion-DPO [68] and D3PO [80] directly optimize the model on pre-collected human preference pairs, bypassing the need for online reward feedback. Building on this paradigm, subsequent works have explored constructing more informative comparison data [85, 32, 29], improved DPO objective [41, 71], and credit assignment over denoising timesteps [43, 90].
Preference learning for video diffusion. Early approaches to preference learning in video diffusion directly borrow techniques from image diffusion, such as direct reward optimization [58, 84, 40, 54, 66] and training loss re-weighting [20], and even relying on image reward models to provide supervision. Recent papers thus focus on developing better video reward models [46, 78, 45]. One strategy aggregates multiple video quality assessment metrics [30, 22] to a final score [46, 87]. However, existing metrics are only effective for short videos [45, 78], limiting their applicability for modern video generators that produce long videos [81, 69, 7]. To address this limitation, LiFT [72], VisionReward [78], and VideoAlign [45] collect a large number of videos from advanced video generators, label human feedback, and fine-tune VLMs to predict preferences. With a powerful video reward model, they apply weighted training [72] or DPO [78, 45] to improve video generation. In contrast to prior works, we focus on DPO for video diffusion using direct human annotations, i.e., without an explicit reward model. Analogous to the verbosity bias observed in language model preference learning, where annotators favor longer outputs [65, 62], we identify a motion bias in video preference labels, where slow-motion videos are often preferred. To mitigate this, we propose a better data pair construction strategy to address this bias via guided video generation.
@@ 3 Method
In this section, we begin by reviewing the background on diffusion models and the standard Direct Preference Optimization (DPO) framework, which we refer to as VanillaDPO (Sec. 3.1). Next, we introduce StructuralDPO, a method that optimizes human preferences on structurally similar video pairs (Sec. 3.2). Finally, we propose DenseDPO, which enables fine-grained human preference alignment along the temporal axis of videos (Sec. 3.3).
@@
Notation. Let $\bm{x}\in\mathbb{R}^{T\times H\times W}$ denote a video sample of length $T$ with spatial dimensions $H\times W$ drawn from the real video distribution $p_{\text{data}}$. We use a diffusion process that transforms real video $\bm{x}$ to Gaussian noise $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$, and reverses this transformation by learning a denoiser $\mathbf{G}_{\bm{\theta}}(\bm{x}_{t},t,\bm{c})$, a neural network parameterized by $\bm{\theta}$, where $\bm{x}_{t}$ is the noisy latent at time $t$ and $\bm{c}$ refers to the auxiliary conditioning variable such as text prompt embeddings or an image.
@@ 3.1 Background: Video Diffusion and DPO
Rectified-flow diffusion models. We follow the rectified flow framework [44, 47], which learns a transport map from the distribution of real videos $\bm{x}\sim p_{\text{data}}$ to the standard normal distribution $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$. The forward diffusion process produces a noisy input $\bm{x}_{t}$ at time $t\in[0,1]$ via a linear interpolation with noise $\bm{\epsilon}$: $\bm{x}_{t}=(1-t)\bm{x}_{0}+t\bm{\epsilon}$. The velocity estimator $\mathbf{G}_{\bm{\theta}}$ is optimized with the following objective:
@@
where $p(t)$ is the distribution of noise levels (following [17], we adopt the logit-normal one).
@@
$l(\bm{x}^{0},\bm{x}^{1})=\begin{cases}+1,&\text{if }\bm{x}^{0}\succ\bm{x}^{1}\,\,\,\,\,(\text{i.e.,\ }\bm{x}^{0}\text{ is preferred over }\bm{x}^{1})\\ -1,&\text{if }\bm{x}^{1}\succ\bm{x}^{0}\,\,\,\,\,(\text{i.e.,\ }\bm{x}^{1}\text{ is preferred over }\bm{x}^{0})\end{cases}$ (2)
@@ 3.2 StructuralDPO: Preference Learning over Structurally Similar Videos
Motion bias in VanillaDPO. In the standard VanillaDPO pipeline, preference pairs are created by independently sampling two videos $(\bm{x}^{0},\bm{x}^{1})$ from different noise seeds under the same conditioning $\bm{c}$ (see Algo. 1), followed by human preference annotation. While this approach works reasonably well for images, its direct extension to videos introduces new issues due to the presence of the new temporal dimension. Independent noises often result in videos with significantly different motion patterns and global layouts (see Fig. 2 (a)). For example, in a typical preference pair, one video may be nearly static but visually clean, while the other contains the desired motion but also introduces artifacts, such as distorted limbs or flickering.
We empirically observe that this is a common bias in generated video data. Indeed, video models often excel at producing high-quality slow-motion clips, while dynamic videos usually contain visible artifacts [8, 86]. Since humans typically perceive clean static videos as more realistic than artifact-prone dynamic ones, this often leads to a preference dataset that systematically over-represents static content. Consequently, a model trained with DPO on this dataset would produce videos with reduced motion. In our preliminary DPO experiments, we observed a substantial drop in dynamic degree using our base model (see Tab. 1 and Tab. 2). The same issue has also been observed in prior works, e.g., Tab.10 of [78] using CogVideoX [81] and Tab. 2 of [46] using VideoCrafter-v2 [11].
@@
The corrupted videos are then denoised from step $n$ to $1$, as outlined in Algo. 2. Here, $\eta$ governs the structural similarity between the two samples. Since early diffusion steps control the global motion [8], this approach preserves the overall dynamics while allowing variations in local details. We compare videos of varying motion strengths generated with different $\eta$ in Fig. 3.
StructuralDPO applies the standard DPO formulation (Eq. 5) on this structurally consistent dataset, which helps to focus the preference on the temporal artifacts and visual inconsistencies, anchoring other dimensions like dynamic degree. Additionally, guided sampling simplifies the generation task, allowing the model to produce artifact-free highly dynamic videos more reliably. Finally, guided sampling reduces data construction costs as it requires fewer sampling steps.
@@
  Input: Denoiser $\mathbf{G}_{\bm{\theta}}$, Input Condition $\bm{c}$, Inference Steps $N$
@@
  Input: Denoiser $\mathbf{G}_{\bm{\theta}}$, Input Condition $\bm{c}$, Inference Steps $N$, Real Video $\bm{x}$, Guidance Level $\eta\in[0,1]$
@@ 3.3 DenseDPO: Rich Temporal Feedback with Segment-Level Preferences
DenseDPO. In VanillaDPO and StructuralDPO, a scalar preference label $l\in\{-1,+1\}$ is obtained for the entire video of length $T$. Instead, DenseDPO annotates preferences on shorter temporal segments. Since guided video generation (Algo. 2) yields structurally similar video pairs, the same time period in both videos has a clear correspondence, making comparison feasible. We show an example in Fig. 2 (b) with the intervals being a single frame: for frame $1$ and $2$, the first video is better, while for frame $3$, the second video is better. Formally, given two videos $(\bm{x}^{0},\bm{x}^{1})$ and the interval length $s$, we can break down videos into $F=\mathrm{ceil}(\frac{T}{s})$ temporal segments of length $s$ by splitting along the time dimension. The resulting video pairs ($(\{\bm{x}^{0}_{f},\bm{x}^{1}_{f}\}^{F}_{f=1})$ are annotated with preferences over each segment, yielding segment-level dense preference labels $\bm{l}\in\{-1,+1\}^{F}$, i.e., $\bm{l}(\bm{x}^{0},\bm{x}^{1})=[l(\bm{x}^{0}_{f},\bm{x}^{1}_{f})]^{F}_{f=1}$. Thus, following Eq. 5, we can formulate the DenseDPO objective as:
@@
In our collected dense preference data, we find that over 60% of video pairs have both winning and losing labels in $\bm{l}$. In regular preference annotation such pairs will either be treated as ties or choose the video with fewer artifacts. In the latter case, this encourages the model to minimize loss on videos with artifacts in Eq. 5, degrading the model performance. In contrast, DenseDPO assigns preference labels more accurately over time, only optimizing models on segments with a clear difference.
Segment preference annotation with VLMs. Another benefit of the DenseDPO is that it allows us to use off-the-shelf VLMs for automatic preference labeling. Prior works point out that existing VLMs struggle at assessing long videos (e.g., 5s) [78, 22], often requiring task-specific fine-tuning and large-scale human annotations to train effective video reward models. Instead, we show that pre-trained VLMs are already capable of processing short clips (e.g., 1s). Given two temporally aligned videos, we feed in pairs of segments into a VLM, and ask it to identify the better one. As we will demonstrate in the experiments (Tab. 3(b)), GPT-o3 [1] achieves high accuracy on segment-level preferences, leading to DPO results competitive with using human preference labels.
@@ 4 Experiments
Our experiments aim to answer the following questions: (i) How does DenseDPO perform against VanillaDPO? (Sec. 4.2) (ii) Can we leverage existing VLMs to produce high-quality preference labels? (Sec. 4.3) (iii) What is the impact of each component in our framework? (Sec. 4.4)
@@ 4.1 Experimental Setup
Baselines. Our pre-trained text-to-video generator is a DiT [53]-based latent flow model. We fine-tune it on the curated high-quality data, termed the SFT baseline. For VanillaDPO, we randomly select 30k text prompts from the curated dataset, generate 2 videos of 5s per prompt with Algo. 1, and ask human labelers to annotate preferences. This leads to around 10k winning-losing pairs after removing ties. For StructuralDPO, we use the same 30k prompts from VanillaDPO, and label human preferences on videos generated using Algo. 2. The guidance level $\eta$ is randomly sampled from $[0.65,0.8]$ to obtain video pairs with similar motion. This again leads to around 10k winning-losing pairs. Both VanillaDPO and StructuralDPO apply the Flow-DPO loss in Eq. 5. Following prior works [78, 45], we set $\beta$ to $500$ and apply LoRA [28] with rank 128 to fine-tune the video model. We train with the AdamW optimizer [48] and a global batch size of 256 for 1000 steps.
@@
Figure 4: Qualitative results. Pre-trained model generates deformed limbs. VanillaDPO fixes it but generates almost static motion. StructuralDPO retains dynamics but produces oversaturated frames. DenseDPO is the only method that generates correct limbs, large dynamics, and high quality visuals.
Figure 5: Human evaluation of DenseDPO vs. StructuralDPO (left) and VanillaDPO (right). TA, VQ, TC, DD stand for text alignment, visual quality, temporal consistency, and dynamic degree.
@@ 4.2 DPO with Human Labels
Tab. 1 and Tab. 2 present the quantitative results on VideoJAM-bench and MotionBench. Fig. 4 shows a qualitative comparison. VanillaDPO greatly improves the pre-trained model and the SFT baseline in all dimensions except dynamic degree due to the motion bias in video preference data. StructuralDPO retains the motion with paired video generation, while compromising the visual quality and text alignment. With the rich temporal feedback, DenseDPO consistently outperforms StructuralDPO. In addition, it matches all aspects of VanillaDPO and scores a significantly higher dynamic degree, despite using only one-third of labeled videos (10k vs. 30k).
Human evaluation. We conduct a user study using all prompts from VideoJAM-bench in Fig. 5. We ask the participants to express their preference when presented with paired samples from our method and each baseline. DenseDPO consistently outperforms StructuralDPO in all dimensions. Compared to VanillaDPO, we achieve significantly higher dynamic degree, and are on par in other aspects.

### v2

@@ Abstract
Direct Preference Optimization (DPO) has recently been applied as a post‑training technique for text-to-video diffusion models. To obtain training data, annotators are asked to provide preferences between two videos generated from independent noise. However, this approach prohibits fine-grained comparisons, and we point out that it biases the annotators towards low-motion clips as they often contain fewer visual artifacts. In this work, we introduce DenseDPO, a method that addresses these shortcomings by making three contributions. First, we create each video pair for DPO by denoising corrupted copies of a ground truth video. This results in aligned pairs with similar motion structures while differing in local details, effectively neutralizing the motion bias. Second, we leverage the resulting temporal alignment to label preferences on short segments rather than entire clips, yielding a denser and more precise learning signal. With only one‑third of the labeled data, DenseDPO greatly improves motion generation over vanilla DPO, while matching it in text alignment, visual quality, and temporal consistency. Finally, we show that DenseDPO unlocks automatic preference annotation using off-the-shelf Vision Language Models (VLMs): GPT accurately predicts segment-level preferences similar to task-specifically fine-tuned video reward models, and DenseDPO trained on these labels achieves performance close to using human labels. Additional results are available at https://snap-research.github.io/DenseDPO/.
@@ 2 Related Work
Preference learning for image diffusion. Inspired by the success of human feedback learning in language modeling [52, 3], similar approaches have been adapted to image generation. One major line of work focuses on training reward models from human preference labels [79, 76, 75, 93, 33], these models can be used as loss functions to optimize generators by direct gradient backpropagation [13, 77, 9, 31, 59, 79] or policy gradients [5, 17]. Another line of work utilizes predicted rewards on training data to re-weight the diffusion loss [38], or trains only on high-scoring samples [15, 14, 39]. However, all these methods require an explicitly trained reward model, and may suffer from the reward hacking issue [79, 13]. In contrast, Diffusion-DPO [70] and D3PO [83] directly optimize the model on pre-collected human preference pairs, bypassing the need for online reward feedback. Building on this paradigm, subsequent works have explored improving comparison data pairs [88, 32, 28, 90], better DPO objectives [41, 73, 29], and credit assignment over denoising timesteps [43, 94].
Preference learning for video diffusion. Early approaches to preference learning in video diffusion directly borrow techniques from image diffusion, such as direct reward optimization [60, 87, 40, 56, 68] and training loss re-weighting [19]. However, they often rely on image reward models [75, 33] to provide supervision. Recent papers thus focus on developing better video reward models [47, 80, 46]. One strategy aggregates multiple video quality assessment metrics [30, 21] to a final score [47, 91]. However, existing metrics are only effective for short videos [46, 80], limiting their applicability for modern video generators that produce long videos [84, 71, 7]. To address this limitation, LiFT [74], VisionReward [80], and VideoAlign [46] collect a large number of videos from advanced video generators, label human feedback, and fine-tune VLMs to predict preferences. With a powerful video reward model, they apply weighted training [74] or DPO [80, 46] to improve video generation. In contrast to prior works, we focus on DPO for video diffusion using direct human annotations, i.e., without an explicit reward model. Analogous to the verbosity bias observed in language model preference learning, where annotators favor longer outputs [67, 64], we identify a motion bias in video preference labels, where slow-motion videos are often preferred. To mitigate this, we propose a better data pair construction strategy to address this bias via guided video generation.
@@ 3 Method
We build upon diffusion models and the standard Direct Preference Optimization (DPO) framework, which we refer to as VanillaDPO (Sec.˜3.1). We discuss the motion bias inherent in using this naïve approach for video generation and introduce StructuralDPO, a method that optimizes human preferences on structurally similar video pairs (Sec.˜3.2). To address the reduction of diversity induced from using structurally similar videos, we propose DenseDPO, which enables fine-grained human preference alignment along the temporal axis of videos (Sec.˜3.3).
@@ 3.1 Background: Video Diffusion and DPO
Rectified-flow diffusion models. Let $\bm{x}\in\mathbb{R}^{T\times H\times W}$ denote a video sample of length $T$ with spatial dimensions $H\times W$. We follow the rectified flow framework [44, 48], which learns a transport map from the standard normal distribution $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$ to the distribution of real videos $\bm{x}\sim p_{\text{data}}$ with a denoiser. The forward diffusion process produces a noisy input $\bm{x}_{t}$ at time $t\in[0,1]$ via a linear interpolation with noise $\bm{\epsilon}$: $\bm{x}_{t}=(1-t)\bm{x}_{0}+t\bm{\epsilon}$. The denoiser $\mathbf{G}_{\bm{\theta}}(\bm{x}_{t},t,\bm{c})$, implemented as a neural network parameterized by $\bm{\theta}$, is trained to reverse this process with the following objective:
@@
where $p(t)$ is the distribution of noise levels (following [16], we adopt the logit-normal one) and $\bm{c}$ refers to the auxiliary conditioning variable such as text embeddings.
@@
[…]
@@ 3.2 StructuralDPO: Preference Learning over Structurally Similar Videos
Motion bias in VanillaDPO. In the standard VanillaDPO pipeline, preference pairs are created by independently sampling two videos $(\bm{x}^{0},\bm{x}^{1})$ from different noise seeds under the same conditioning $\bm{c}$ (see Algo.˜1), followed by human preference annotation. While this approach works reasonably well for images, its direct extension to videos introduces new issues due to the presence of the new temporal dimension. Independent noises often result in videos with significantly different motion patterns and global layouts (see Fig.˜2 (a)). For example, in a typical preference pair, one video may be nearly static but visually clean, while the other contains the desired motion but also introduces artifacts, such as distorted limbs or flickering.
We empirically observe that this is a common bias in generated video data. Indeed, video models often excel at producing high-quality slow-motion clips, while dynamic videos usually contain visible artifacts [8, 89]. Since humans typically perceive clean static videos as more realistic than artifact-prone dynamic ones, this often leads to a preference dataset that systematically over-represents static content. Consequently, a model trained with DPO on this dataset would produce videos with reduced motion. In our preliminary DPO experiments, we observed a substantial drop in dynamic degree using our base model (see Tab.˜1 and Tab.˜2). The same issue has also been observed in prior works, e.g., Tab.10 of [80] using CogVideoX [84] and Tab. 2 of [47] using VideoCrafter-v2 [11].
@@
The corrupted videos are then denoised from step $n$ to $1$, as outlined in Algo.˜2. Here, $\eta$ governs the structural similarity between the two samples. Since early diffusion steps control the global motion [8], this approach preserves the overall dynamics while allowing variations in local details. We compare videos of varying motion strengths generated with different $\eta$ in Fig.˜3.
StructuralDPO applies the standard DPO formulation (Eq.˜5) on this structurally consistent dataset, which helps to focus the preference on the temporal artifacts and visual inconsistencies, anchoring other dimensions like dynamic degree. Additionally, guided sampling simplifies the generation task, allowing the model to produce artifact-free highly dynamic videos more reliably. Finally, guided sampling reduces data construction costs as it requires fewer sampling steps.
@@
Input: Denoiser $\mathbf{G}_{\bm{\theta}}$, Input Condition $\bm{c}$, Inference Steps NN
@@
Input: Denoiser $\mathbf{G}_{\bm{\theta}}$, Input Condition $\bm{c}$, Inference Steps NN, Real Video $\bm{x}$, Guidance Level $\eta\in[0,1]$
@@ 3.3 DenseDPO: Rich Temporal Feedback with Segment-Level Preferences
DenseDPO. In VanillaDPO and StructuralDPO, a scalar preference label $l\in\{-1,+1\}$ is obtained for the entire video of length $T$. Instead, DenseDPO annotates preferences on shorter temporal segments. Since guided video generation (Algo.˜2) yields structurally similar video pairs, the same time period in both videos has a clear correspondence, making comparison feasible. We show an example in Fig.˜2 (b) with the intervals being a single frame: for frame $1$ and $2$, the first video is better, while for frame $3$, the second video is better. Formally, given two videos $(\bm{x}^{0},\bm{x}^{1})$ and the interval length $s$, we can break down videos into $F=\mathrm{ceil}(\frac{T}{s})$ temporal segments of length $s$ by splitting along the time dimension. The resulting video pairs $(\{\bm{x}^{0}_{f},\bm{x}^{1}_{f}\}^{F}_{f=1})$ are annotated with preferences over each segment, yielding segment-level dense preference labels $\bm{l}\in\{-1,+1\}^{F}$, i.e., $\bm{l}(\bm{x}^{0},\bm{x}^{1})=[l(\bm{x}^{0}_{f},\bm{x}^{1}_{f})]^{F}_{f=1}$. Thus, following Eq.˜5, we can formulate the DenseDPO objective as:
@@
In our collected dense preference data, we find that over 60% of video pairs have both winning and losing labels in $\bm{l}$. In regular preference annotation such pairs will either be treated as ties or choose the video with fewer artifacts. In the latter case, this encourages the model to minimize loss on videos with artifacts in Eq.˜5, degrading the model performance. In contrast, DenseDPO assigns preference labels more accurately over time, only optimizing models on segments with a clear difference.
Segment preference annotation with VLMs. Another benefit of the DenseDPO is that it allows us to use off-the-shelf VLMs for automatic preference labeling. Prior works point out that existing VLMs struggle at assessing long videos (e.g., 5s) [80, 21], often requiring task-specific fine-tuning and large-scale human annotations to train effective video reward models. Instead, we show that pre-trained VLMs are already capable of processing short clips (e.g., 1s). Given two temporally aligned videos, we feed in pairs of segments into a VLM, and ask it to identify the better one. As we will demonstrate in the experiments (Tab.˜3(b)), GPT-o3 [1] achieves high accuracy on segment-level preferences, leading to DPO results competitive with using human preference labels.
@@ 4 Experiments
Our experiments aim to answer the following questions: (i) How does DenseDPO perform against VanillaDPO? (Sec.˜4.2) (ii) Can we leverage existing VLMs to produce high-quality preference labels? (Sec.˜4.3) (iii) What is the impact of each component in our framework? (Sec.˜4.4)
@@ 4.1 Experimental Setup
Baselines. Our pre-trained text-to-video generator is a DiT [55]-based latent flow model. We fine-tune it on the curated high-quality data, termed the SFT baseline. For VanillaDPO, we randomly select 30k text prompts from the curated dataset, generate 2 videos of 5s per prompt with Algo.˜1, and ask human labelers to annotate preferences. This leads to around 10k winning-losing pairs after removing ties. For StructuralDPO, we use the same 30k prompts from VanillaDPO, and label human preferences on videos generated using Algo.˜2. The guidance level $\eta$ is randomly sampled from $[0.65,0.8]$ to obtain video pairs with similar motion. This again leads to around 10k winning-losing pairs. Both VanillaDPO and StructuralDPO apply the Flow-DPO loss in Eq.˜5. Following prior works [80, 46], we set $\beta$ to $500$ and apply LoRA [27] with rank 128 to fine-tune the video model. We train with the AdamW optimizer [49] and a global batch size of 256 for 1000 steps.
@@
Figure 4: Qualitative results. Pre-trained model generates deformed limbs. VanillaDPO fixes it but generates almost static motion. StructuralDPO retains dynamics but produces oversaturated frames. DenseDPO is the only method that generates correct limbs, large dynamics, and high quality visuals. Please check out our project page for video results of baselines and our methods.
Figure 5: Human evaluation of DenseDPO vs. StructuralDPO (left) and VanillaDPO (right). TA, VQ, TC, $D$ stand for text alignment, visual quality, temporal consistency, and dynamic degree.
@@ 4.2 DPO with Human Labels
Tab.˜1 and Tab.˜2 present the quantitative results on VideoJAM-bench and MotionBench. Fig.˜4 shows a qualitative comparison. VanillaDPO greatly improves the pre-trained model and the SFT baseline in all dimensions except dynamic degree due to the motion bias in video preference data. StructuralDPO retains the motion with paired video generation, while compromising the visual quality and text alignment. With the rich temporal feedback, DenseDPO consistently outperforms StructuralDPO. In addition, it matches all aspects of VanillaDPO and scores a significantly higher dynamic degree, despite using only one-third of labeled videos (10k vs. 30k).
Please refer to Appendix C.1 for comparison with more baselines including online RL-based methods. For more qualitative results, please check out Appendix C.2 and our project page for video results.
Human evaluation. We conduct a user study using all prompts from VideoJAM-bench in Fig.˜5. We ask the participants to express their preference when presented with paired samples from our method and each baseline. DenseDPO consistently outperforms StructuralDPO in all dimensions. Compared to VanillaDPO, we achieve significantly higher dynamic degree, and are on par in other aspects. Please refer to Appendix C.2 for more user study results.